OpenStudy (anonymous):

What is 15-10+2/5*-2

OpenStudy (anonymous):

At the risk of this being a trick question, is it \[\Large 15 -10 +\frac25\times-2\]?

OpenStudy (anonymous):

I think so..

OpenStudy (anonymous):

OR \[\Large 15 -10+2\div5\times-2\]

OpenStudy (anonymous):

Here first multiply -2 with 2/5..

OpenStudy (anonymous):

It might not actually matter either way :) MDAS, anyone? ^.^

OpenStudy (anonymous):

\[\frac{2}{5} = ??\]

OpenStudy (anonymous):

I really didn't like how MDAS is so misleading, it's actually more like (MD) (AS) Perform the multiplications or divisions first, whichever is leftmost And then perform the additions and subtractions, whichever is leftmost :D

OpenStudy (anonymous):

So, perform the division first... \[\Large 15-10+\color{blue}{2\div5}\times-2\]

OpenStudy (anonymous):

Either you will do multiply 2/5 with -2 or you will first divide and then multiply it with -2, you will get the same in both the cases..

OpenStudy (anonymous):

So, let's hop to it, though I'd rather not continue unless the asker shows him/herself, first :)

OpenStudy (anonymous):

@flying.blast your involvement is missing..

OpenStudy (anonymous):

I am so sorry. I was taking a shower and i just finished. is is 2 divided by 5

OpenStudy (anonymous):

so i divide first? 2/5=0.4 So what next?

OpenStudy (anonymous):

Okay, so we now have \[\Large 15-10+0.4\times2\]What is the next operation to perform?

OpenStudy (anonymous):

uhhh. is it subtracting?

OpenStudy (anonymous):

15-10=5

OpenStudy (anonymous):

No. MDAS

OpenStudy (anonymous):

What is MDAS?

OpenStudy (anonymous):

multiplication division addition subtraction

OpenStudy (anonymous):

OHHHHHHHHHH! So it is 5*-2=-10 and then 2/-10=-0.2 and then 10+(-0.2)=9.8 and 15-9.8=5.2 Am I right?

OpenStudy (anonymous):

Take it slow :) \[\Large 15-10+\color{blue}{0.4\times2}\]

OpenStudy (anonymous):

oh. 15-10+0.8

OpenStudy (anonymous):

And from there, you can pretty much do the operations in any order you want :D

OpenStudy (anonymous):

Or not. Do the subtraction first. As it is left-most

OpenStudy (anonymous):

15-10= 5 and then 5+0.8=5.8 Is that right?

OpenStudy (anonymous):

That's done :)

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

Thank you SO MUCH!!!

OpenStudy (anonymous):

NO!!!

OpenStudy (anonymous):

I forgot a minus-sign \[\Large 15-10+\color{blue}{0.4\times-2}\]

OpenStudy (anonymous):

ohhhhhh. So it is 4.2

OpenStudy (anonymous):

Well THNX AGAIN!!